OpenStudy (anonymous):

v = s^2 + 1/2sh, solve for h

OpenStudy (lgbasallote):

\[V = s^2 + \frac 12 sh\] subtract s^2 from both sides \[\implies V - s^2 = \frac 12 sh\] multiply 2 to both sides \[\implies 2(V-s^2) = sh\] divide both sides by s \[\implies \frac{2(V - s^2)}{s} = h\] does that help?

OpenStudy (anonymous):

@lgbasallote Yes very much!!! Thank you so much. :)

OpenStudy (lgbasallote):

welcome